OpenStudy (student852):

Quinn invests $7500 at a bank that offers a 5.5% annual interest rate compounded daily. If t represents time in years, which expression represents the approximate amount of money in the account?

OpenStudy (student852):

A.7500(1.00015)^t B. 7500(1.00015)^3615t C. 7500(1.55)^t D. 7500(1.55)^365t

OpenStudy (student852):

pleaseeeee help

OpenStudy (kevin):

i = 5.5% = 0.055 Mn = Mo (1 + i)^n Mn = 7500 (1 + 0.055)^t Mn = 7500 (1.055)^t

OpenStudy (student852):

thanks!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!